OpenStudy (anonymous):

Assume a 115 pound person burns 36 calories after 12 minutes of walking and 54 calories after 18 minutes of walking. Part 1 (2 points): Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. Part 3 (2 points): How many calories will this person burn after 35 minutes of walking? Using complete sentences, explain how the equation, slope, or graph can help to predict calories

OpenStudy (anonymous):

(12,36) and (18,54) m=y2-y1/x2-x1 54-36/18-12=18/6=3 3 or 3/1 is the slope. y=3x+b 36=3(12)+b 36=36+b -36-36 0=b check: 54=3(18)+b 54=54+b -54-54 0=b y=3x+0 35=3x+0 35=3x -- --- 3 3 11.667=x (11.667,35)

OpenStudy (anonymous):

which is part 1 2 and 3?

OpenStudy (anonymous):

Part 1 (12,36) and (18,54) m=y2-y1/x2-x1 54-36/18-12=18/6=3 3 or 3/1 is the slope. y=3x+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Part 2 36=3(12)+b 36=36+b -36-36 0=b check: 54=3(18)+b 54=54+b -54-54 0=b y=3x+0

OpenStudy (anonymous):

Part 3 35=3x+0 35=3x -- --- 3 3 11.667=x (11.667,35)